If the mean of a set of data is 24, and 18.40 has a z

-

score of

1.40, then the

variance must be:



1
Expert's answer
2021-08-10T10:12:08-0400

Solution:

Given, "\mu=24,\bar X=18.40,z=-1.40,\sigma=?"

"z=\dfrac{\bar X-\mu}{\sigma}
\\ \Rightarrow-1.40=\dfrac{18.40-24}{\sigma}
\\ \Rightarrow \sigma=\dfrac{-5.6}{-1.40}=4"